OpenStudy (anonymous):

square root of 3 equals 45 over x

OpenStudy (lgbasallote):

\[\sqrt 3 = \frac{45}x\] cross multiply... \[\implies x \sqrt 3 = 45\] do you know how to solve for x now?

OpenStudy (anonymous):

maybe.. I got 45 square root 3 over 3?

OpenStudy (anonymous):